OpenStudy (anonymous):

find the limit lim x approaches 13 for (13-x)/(x^2-169)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 13} \frac{(13-x)}{(x^2-169)} \] \[=\lim_{x \rightarrow 13} \frac{(13-x)}{(x^2-13^2)} \] \[=\lim_{x \rightarrow 13} \frac{(13-x)}{(x+13)(x-13)} \] \[=\lim_{x \rightarrow 13} \frac{(13-x)}{-(x+13)(13-x)} =\lim_{x \rightarrow 13} \frac{-1}{(x+13)} \] Applying limit we find \[ = \frac{-1}{(13+13)} = \frac{-1}{26} \]

OpenStudy (anonymous):

@nlpowell

OpenStudy (anonymous):

Thank you.